# Untold Stories: Unveiling the Scarce Contributions of UX Professionals to Usability Issue Discussions of Open Source Software Projects

### v1

@@ 1. Introduction
RQ2: How do UX professionals follow up with the usability issues they posted? After characterizing the issues reported by the UX professionals, we investigated how they continued to contribute to resolving those issues they posted. Concretely, we analyzed their behavior after posting usability issues, when following up on those issues in comments. Through this analysis, we found that UX professionals engaged in about 1/3 of their issues as commenters. An inductive coding on the purpose of their follow-up comments indicated that their main goals were to share ideas/opinions/experiences and provide supplementary details so that their feedback or suggestions could be better understood, taken seriously, and addressed by the OSS developers.
@@ 3.1.3. Argument Structure
3.2. RQ2: How Do UX Professionals Follow Up With the Usability Issues They Posted?
@@ 3.2.2. UX Professionals’ Purposes for Following up on Their Usability Issues
Awesome, my real secret was I was waiting for @benogle’s thoughts to go forward :) Next matter to discus regarding these tabs is… Separate tabs from UI & Syntax Styles: I’ve been noticing in my discovery that many themes have dark tabs/syntax and a light theme around the web. Others may prefer lighter UITabs with a dark UI. One thought would be to make them separate from the UI and the Syntax. For myself I’d match them up ’usually’ with the syntax and not the UI. Anyone with strong thoughts on this?
@@
Giving extra information/example: The participants of comments contributed additional details, examples, or context to enrich the discussion. Their aim was to provide supplementary information or examples for a better understanding of a discussed topic. Examples of this type of purpose are VSCode UX professionals in issue #4331:
This will also be relevant to #4100” and UX professional male in Jupyter Lab of #6615: “@ellisonbg the folder icon is rendering in the correct alignment on my JupyterLab. But I haven’t updated anything since yesterday; maybe it’s a regression? I’m on Chrome Version 74.0.3729.169 (Official Build) (64-bit) and the latest OS.
@@
Expressing enthusiasm. The commenters sometimes exhibited excitement, positive feedback, or encouragement regarding the topic or solution being discussed. Their goal in the comment was to express support or appreciation. One example of this is from Atom’s UX professionals in issue #840: “Look forward to updating to check this out.
@@ 4.1.1. Scarcity of UX Professionals’ Engagement in OSS Usability Issue Discussions
Our results yielded crucial evidence that UX professionals’ involvement in the OSS issue discussions is very scarce; i.e., only four individuals were identified as UX professionals from the 224 usability issue posters. This limited presence raises concerns about the potential impact on UX quality and overall user satisfaction in OSS projects. We recognized that one possible reason behind their limited involvement in the issue tracking system is that they may conduct and discuss their work related to usability, such as user studies and UX design, outside of these systems. However, we argue that because the central development decisions, project management, and collaboration for OSS take place within ITSs, the limited participation of UX professionals on these platforms creates a potential disconnection between the design considerations that they offer and the actual development workflow. This disconnect also is problematic as it may lead to a lack of alignment between design goals, overall project objectives, and lack of user satisfaction. Thus, while we recognize that UX and usability-related tasks go beyond issue discussions, understanding the implications of limited contribution by UX professionals in these systems is essential for addressing challenges and fostering more effective collaboration in OSS development.

### v2

@@ 1. Introduction
RQ2: How do UX professionals follow up on the usability issues they posted? After characterizing the issues reported by the UX professionals, we investigated how they continued to contribute to resolving those issues they posted. Concretely, we analyzed their behavior after posting usability issues, when following up on those issues in comments. Through this analysis, we found that UX professionals engaged in about 1/3 of their issues as commenters. An inductive coding on the purpose of their follow-up comments indicated that their main goals were to share ideas/opinions/experiences and provide supplementary details so that their feedback or suggestions could be better understood, taken seriously, and addressed by the OSS developers.
@@ 3.1.3. Argument Structure
3.2. RQ2: How Do UX Professionals Follow Up on the Usability Issues They Posted?
@@ 3.2.2. UX Professionals’ Purposes for Following up on Their Usability Issues
Awesome, my real secret was I was waiting for @benogle’s thoughts to go forward :) Next matter to discuss regarding these tabs is… Separate tabs from UI & Syntax Styles: I’ve been noticing in my discovery that many themes have dark tabs/syntax and a light theme around the web. Others may prefer lighter UITabs with a dark UI. One thought would be to make them separate from the UI and the Syntax. For myself I’d match them up ‘usually’ with the syntax and not the UI. Anyone with strong thoughts on this?
@@
Giving extra information/example: The participants of comments contributed additional details, examples, or context to enrich the discussion. Their aim was to provide supplementary information or examples for a better understanding of a discussed topic. Examples of this type of purpose are VSCode UX professionals in issue #4331: “This will also be relevant to #4100” and the male UX professional in Jupyter Lab of #6615:
@ellisonbg the folder icon is rendering in the correct alignment on my JupyterLab. But I haven’t updated anything since yesterday; maybe it’s a regression? I’m on Chrome Version 74.0.3729.169 (Official Build) (64-bit) and the latest OS.
@@
Expressing enthusiasm: The commenters sometimes exhibited excitement, positive feedback, or encouragement regarding the topic or solution being discussed. Their goal in the comment was to express support or appreciation. One example of this is from Atom’s UX professionals in issue #840: “Look forward to updating to check this out.
@@ 4.1.1. Scarcity of UX Professionals’ Engagement in OSS Usability Issue Discussions
Our results yielded crucial evidence that UX professionals’ involvement in the OSS issue discussions is very scarce; i.e., only four individuals were identified as UX professionals from the 224 usability issue posters. This limited presence raises concerns about the potential impact on UX quality and overall user satisfaction in OSS projects. We recognized that one possible reason behind their limited involvement in the issue tracking system is that they may conduct and discuss their work related to usability, such as user studies and UX design, outside of these systems. However, we argue that because the central development decisions, project management, and collaboration for OSS take place within ITSs, the limited participation of UX professionals on these platforms creates a potential disconnection between the design considerations that they offer and the actual development workflow. This disconnection is also problematic as it may lead to a lack of alignment between design goals, overall project objectives, and lack of user satisfaction. Thus, while we recognize that UX and usability-related tasks go beyond issue discussions, understanding the implications of limited contribution by UX professionals in these systems is essential for addressing challenges and fostering more effective collaboration in OSS development.